# Optimal Rates for Generalization of Gradient Descent for Deep ReLU Classification

### v1

@@ 1 Introduction
Deep neural networks trained through first-order optimization methods have achieved remarkable empirical success in diverse domains (Krizhevsky et al., 2012). Despite their widespread adoption, a rigorous theoretical understanding of their optimization dynamics and generalization behavior remains incomplete, particularly for ReLU networks. The inherent challenges arise from the non-smoothness and non-convexity of the loss landscape induced by ReLU activations and network architectures, which complicate the classical analysis. Intriguingly, empirical evidence demonstrates that over-parameterized models often achieve zero training error but still generalize well even in the absence of explicit regularization (Zhang et al., 2016). This phenomenon has spurred significant theoretical research to understand its underlying mechanisms.
@@
A prominent line of research uses the neural tangent kernel (NTK) framework to analyze neural network training (Jacot et al., 2018). In the infinite-width limit, gradient descent (GD) dynamics can be characterized by functions in the NTK’s reproducing kernel Hilbert space (RKHS), with convergence guarantees established for both shallow and deep networks (Du et al., 2019b, a). These results demonstrate that GD converges to global minima within a local neighborhood of initialization, provided that the network width is sufficiently large. In particular, the appealing work (Arora et al., 2019a) showed that, if the network width $m=\widetilde{O}(\mathrm{poly}(n,1/\lambda_{0},L))$, then the generalization error is of the order $L\sqrt{\frac{\mathbf{y}^{\top}(\mathbf{H}^{\infty})^{-1}\mathbf{y}}{n}}$, where $\mathbf{H}^{\infty}$ denotes the NTK gram matrix over the training data and $\lambda_{0}=\lambda_{\min}(\mathbf{H}^{\infty})$. However, the assumption $\lambda_{0}>0$ is a strong assumption because it tends to zero if the size of the training data tends to infinity as shown by Su and Yang (2019). Ji and Telgarsky (2020) achieved logarithmic width requirements for NTK-separable data with a margin $\gamma$. They derived the risk bound of order $1/(\gamma^{2}\sqrt{n})$ for two-layer ReLU networks using Rademacher complexity. The recent work Chen et al. (2021) extended their results from shallow to deep neural networks, the authors considered the NTK feature learning and proved the bound of $\widetilde{O}\left(\frac{e^{O(L)}}{\gamma^{2}}\sqrt{\frac{m}{n}}\right)$. However, all the above bounds explicit $\frac{1}{\sqrt{n}}$ dependence on the sample size $n$, which is suboptimal.
@@
Complementing the NTK framework, another line of research employs algorithmic stability to analyze neural networks. In particular, Liu et al. (2020) demonstrated that the Hessian spectral norm scales with width $m$ as $\widetilde{O}\left(\frac{1}{\sqrt{m}}\right)$, providing the theoretical basis for subsequent studies on generalization in overparameterized networks (Richards and Kuzborskij, 2021; Lei et al., 2022; Taheri and Thrampoulidis, 2024; Taheri et al., 2025). The work (Taheri and Thrampoulidis, 2024) achieves the bound of $\widetilde{O}(1/n\gamma^{2})$ for shallow neural networks, which is almost optimal, as illustrated by Shamir (2021); Schliserman and Koren (2023). More recently, Taheri et al. (2025) extended this line of work to deep networks, obtaining a generalization bound of $\widetilde{O}(e^{O(L)}/(n\gamma^{2}))$. However, this approach often assume smooth activation functions and can not apply to the non-smooth ReLU networks. In summary, these works either consider smooth neural networks or develop generalization bounds with exponential dependency on $L$. These limitations motivate two fundamental questions:
@@
Ji and Telgarsky ReLU $\widetilde{\Omega}\left(\frac{1}{\gamma^{8}}\right)$ $\widetilde{O}\left(\frac{1}{\gamma^{2}T}\right)$ $\widetilde{O}\left(\frac{1}{\gamma^{2}\sqrt{n}}\right)$ Shallow
Chen et al. ReLU $\widetilde{\Omega}\left(\frac{L^{22}}{\gamma^{8}}\right)$ $\widetilde{O}\left(\frac{1}{\gamma^{2}T}\right)$ $\widetilde{O}\left(\frac{e^{O(L)}}{\gamma^{2}}\sqrt{\frac{m}{n}}\right)$ Deep
Taheri et al. Smooth $\widetilde{\Omega}\left(\frac{1}{\gamma^{6L+4}}\right)$ $\widetilde{O}\left(\frac{1}{\gamma^{2}T}\right)$ $\widetilde{O}\left(\frac{e^{O(L)}}{\gamma^{2}n}\right)$ Deep
@@
    We prove that gradient decent with step size $\eta$ and $T$ iterations achieves the convergence rate of $F_{S}(\overline{\mathbf{W}})/(\eta T)$, where $F_{S}(\overline{\mathbf{W}})=3\eta T\mathcal{L}_{S}(\overline{\mathbf{W}})+\|\overline{\mathbf{W}}-\mathbf{W}(0)\|_{F}^{2}$, $\overline{\mathbf{W}}$ is a reference model, $\mathbf{W}(0)$ is the initialization point and $\mathcal{L}_{S}(\cdot)$ is the training error. This indicates that all iterates remain within a neighborhood of $\overline{\mathbf{W}}$. By refining the analysis of ReLU activation patterns, we reduce the overparameterization requirement by a factor of $L^{6}$ as compared to Chen et al. (2021) (see Remark 1 ).
@@
    We establish a population risk bound of $\widetilde{O}(L^{4}F(\overline{\mathbf{W}})/n)$, where $F(\overline{\mathbf{W}})$ extends the empirical counterpart $F_{S}(\overline{\mathbf{W}})$ to the population setting. Our analysis introduces two key technical contributions. First, we derive sharper Rademacher complexity bounds for the hypothesis class encompassing all gradient descent iterates. A central innovation is our representation of the complexity via products of sparse matrices, whose norms are tightly controlled using optimization-informed estimates (see Remark 2). Second, by leveraging the covering number techniques, we prove that ReLU networks remain $\widetilde{O}(L^{2})$-Lipschitz continuous in a neighborhood around the initialization—a substantial improvement over previous exponential bounds (Xu and Zhu, 2024; Taheri et al., 2025) (see Remark 3).
@@
    For NTK separable data with a margin $\gamma$, we show that neural networks with $m=\widetilde{O}(L^{16}\log({n/\delta})(\log n)^{8}/\gamma^{8})$ can achieve $\widetilde{O}(L^{4}(1+\gamma L^{2})/(\gamma^{2}n))$ risk. This improves on Chen et al. (2021)’s $\widetilde{O}(e^{O(L)}\gamma^{-2}\sqrt{m/n})$ by simultaneously (a) removing the exponential depth dependence, (b) eliminating the $\sqrt{m}$ width factor and (c) achieving the optimal dependence on sample complexity (see Table 1 for a comparison with existing work).
@@ 2.1 Optimization
The foundational work of Jacot et al. (2018) introduced the Neural Tangent Kernel (NTK) framework, demonstrating that in the infinite-width limit, neural networks behave as linear models with a fixed tangent kernel (Liu et al., 2020; Lee et al., 2019). This lazy training regime (Chizat et al., 2019), where parameters remain close to initialization, enables gradient descent to converge to global optima near initialization (Du et al., 2019a; Arora et al., 2019a). These analyses showed that the training dynamics can be governed by the NTK Gram matrix, which leads to substantial overparameterization ($m\gtrsim n^{6}/\lambda_{0}^{4}$). This was later improved by Oymak and Soltanolkotabi (2020). They showed that if the square-root of the number of the network parameters exceeds the size of the training data, randomly initialized gradient descent converges at a geometric rate to a nearby global optima. The work (Ji and Telgarsky (2020)) achieved polylogarithmic width requirements for logistic loss by leveraging the 1-homogeneity of two-layer ReLU networks. However, it should be mentioned that this special property does not hold for deep networks. The NTK framework was extended to deep architectures by Arora et al. (2019b) for CNNs and by Du et al. (2019b) for ResNets using the last-layer NTK. Xu and Zhu (2024) pointed out that such a characterization is loose, only capturing the contribution from the last layer. They further gave the uniform convergence of all layers as $m\to\infty$ and convergence guarantee for stochastic gradient descent (SGD) in streaming data setting. Allen-Zhu et al. (2019) showed that the optimization landscape is almost-convex and semi-smooth, based on which they proved that SGD can find global minima. Cao and Gu (2019) introduced the neural tangent random feature and showed the convergence of SGD under the overparameterized assumption $m\gtrsim n^{7}$.
@@ 2.2 Generalization
The NTK framework has yielded generalization bounds scaling as $\sqrt{\mathbf{y}^{\top}(\mathbf{H}^{\infty})^{-1}\mathbf{y}/n}$ (Arora et al., 2019a; Cao and Gu, 2019). This data-dependent complexity measure helps to distinguish between random labels and true labels. Li and Liang (2018) showed that SGD trained networks can achieve small test error on specific structured data. A very popular approach to studying the generalization of neural networks is via the uniform convergence, which analyzes generalization gaps in a hypothesis space using tools such as Rademacher complexity or covering numbers (Neyshabur et al., 2015; Bartlett et al., 2017; Golowich et al., 2018; Liu et al., 2024). However, this could lead to vacuous generalization bound in some cases (Nagarajan and Kolter, 2019). Moreover, these bounds typically exhibit exponential dependence on depth $L$, thus often leading to loose bounds (Chen et al., 2021). This capacity-based method usually results in the generalization rate of the order $\widetilde{O}(1/\sqrt{n})$. Recent work has also exploited stability arguments for generalization guarantees (Richards and Kuzborskij, 2021; Lei et al., 2022; Taheri and Thrampoulidis, 2024; Deora et al., 2023; Taheri et al., 2025). The main idea of algorithmic stability is to study how the perturbation of training samples would affect the output of an algorithm (Rogers and Wagner, 1978). The connection to generalization bound was established in Bousquet and Elisseeff (2002). Hardt et al. (2016) gave the stability analysis of SGD for Lipschitz, smooth and convex problems. Lei and Ying (2020) further studied SGD under much wilder assumptions. Liu et al. (2020) identified weak convexity of neural networks, enabling stability analyses with polynomial width requirements for quadratic loss (Richards and Kuzborskij, 2021; Lei et al., 2022). Moreover, Taheri and Thrampoulidis (2024); Taheri et al. (2025) obtained generalization bounds of order $\widetilde{O}(1/n)$ by using a generalized local quasi-convexity property for sufficiently parameterized networks. However, these methods depend on smooth activations, and whether similar or even better bound can be established for deep ReLU networks is still unknown.
@@ 3 Preliminaries
where we use logistic loss $\ell(z):=\log(1+\exp(-z))$ throughout this paper. We further assume the following symmetric initialization (Nitanda and Suzuki, 2020; Kuzborskij and Szepesvári, 2023; Xu and Zhu, 2024):
@@
We remark that this initialization is for theoretical simplicity, using general initialization techniques will not affect the main results. We fix the output weights $\{a_{r}\}$ and use Gradient Descent (GD) to train the weight matrix $\mathbf{W}$ (Ji and Telgarsky, 2020; Arora et al., 2019a; Zou et al., 2018).
@@
Note that in each layer we employ $\sqrt{{2}/{m}}$ as the regularization factor instead of the conventional $\sqrt{{1}/{m}}$ (Ji and Telgarsky, 2020), which is due to $\mathbb{E}_{x\sim\mathcal{N}(0,1)}\sigma^{2}(x)={1}/{2}$ for our activation function $\sigma(\cdot)$. This scaling matches both the theoretical framework of Du et al. (2019a) and the initialization scheme of (He et al., 2015) (where weights $\mathbf{w}_{r}^{l}\sim\mathcal{N}(0,{2}/{m})$). As will be shown later (Appendix A), this regularization ensures stable gradient propagation and maintains consistent variance across layers.
The following assumption is standard in the literature (Cao and Gu, 2019; Ji and Telgarsky, 2020; Chen et al., 2021).
@@
where the first term captures the generalization error, quantifying the network’s performance on unseen data. The second term represents the optimization error, which reflects GD’s ability to find global minima. We will use tools in the optimization theory to study the optimization error (Ji and Telgarsky, 2020; Schliserman and Koren, 2022), and Rademacher complexity to control the generalization error (Mohri, 2018).
@@ 4.1 Optimization Analysis
Our theorem shows that the convergence rate is bounded by the optimization error of a reference model, implying that any low-loss reference point guarantees good convergence. While prior works relied on NTK-induced solutions (Richards and Kuzborskij, 2021; Arora et al., 2019a), we prove that there exists a reference model near initialization under the milder Assumption 3. Furthermore, our analysis implies that all training iterates remain within a neighborhood of the reference point, and thus near initialization, aligned with NTK-based observations but without studying the kernel or the corresponding Gram matrix (Du et al., 2019a, b).
@@
Telescoping gives the theorem. Chen et al. (2021) introduce the following neural tangent random feature (NTRF) function class:

### v2

@@ 1 Introduction
Deep neural networks trained through first-order optimization methods have achieved remarkable empirical success in diverse domains (NIPS2012_c399862d). Despite their widespread adoption, a rigorous theoretical understanding of their optimization dynamics and generalization behavior remains incomplete, particularly for ReLU networks. The inherent challenges arise from the non-smoothness and non-convexity of the loss landscape induced by ReLU activations and network architectures, which complicate the classical analysis. Intriguingly, empirical evidence demonstrates that over-parameterized models often achieve zero training error but still generalize well even in the absence of explicit regularization (zhang2016understanding). This phenomenon has spurred significant theoretical research to understand its underlying mechanisms.
@@
A prominent line of research uses the neural tangent kernel (NTK) framework to analyze neural network training (jacot2018neural). In the infinite-width limit, gradient descent (GD) dynamics can be characterized by functions in the NTK’s reproducing kernel Hilbert space (RKHS), with convergence guarantees established for both shallow and deep networks (du2019gradientdescentprovablyoptimizes; pmlr-v97-du19c). These results demonstrate that GD converges to global minima within a local neighborhood of initialization, provided that the network width is sufficiently large. In particular, the appealing work (pmlr-v97-arora19a) showed that, if the network width $m=\widetilde{O}(\mathrm{poly}(n,1/\lambda_{0},L))$, then the generalization error is of the order $L\sqrt{\frac{\mathbf{y}^{\top}(\mathbf{H}^{\infty})^{-1}\mathbf{y}}{n}}$, where $\mathbf{H}^{\infty}$ denotes the NTK gram matrix over the training data and $\lambda_{0}=\lambda_{\min}(\mathbf{H}^{\infty})$. However, the assumption $\lambda_{0}>0$ is a strong assumption because it tends to zero if the size of the training data tends to infinity as shown by su2019learning. ji2020polylogarithmicwidthsufficesgradient achieved logarithmic width requirements for NTK-separable data with a margin $\gamma$. They derived the risk bound of order $1/(\gamma^{2}\sqrt{n})$ for two-layer ReLU networks using Rademacher complexity. The recent work chen2021overparameterizationsufficientlearndeep extended their results from shallow to deep neural networks, the authors considered the NTK feature learning and proved the bound of $\widetilde{O}\left(\frac{e^{O(L)}}{\gamma}\sqrt{\frac{m}{n}}\right)$. Recently, lei2024optimization derived the bound of $1/(\gamma^{2}n)$ for two-layer ReLU networks. However, all the above bounds explicit suboptimal $\frac{1}{\sqrt{n}}$ dependence on the sample size $n$ or only focus on shallow networks.
@@
Complementing the NTK framework, another line of research employs algorithmic stability to analyze neural networks. In particular, NEURIPS2020_b7ae8fec demonstrated that the Hessian spectral norm scales with width $m$ as $\widetilde{O}\left(\frac{1}{\sqrt{m}}\right)$, providing the theoretical basis for subsequent studies on generalization in overparameterized networks (richards2021stability; NEURIPS2022_fb8fe6b7; JMLR:v25:23-0422; taheri2025sharper). The work (JMLR:v25:23-0422) achieves the bound of $\widetilde{O}(1/n\gamma^{2})$ for shallow neural networks, which is almost optimal, as illustrated by JMLR:v22:20-997; NEURIPS2023_d8ca28a3. More recently, taheri2025sharper extended this line of work to deep networks, obtaining a generalization bound of $\widetilde{O}(e^{O(L)}/(n\gamma^{2}))$. However, this approach often assume smooth activation functions and can not apply to the non-smooth ReLU networks. In summary, these works either consider smooth neural networks or develop generalization bounds with exponential dependency on $L$. These limitations motivate two fundamental questions:
@@
ji2020polylogarithmicwidthsufficesgradient ReLU $\widetilde{\Omega}\left(\frac{1}{\gamma^{8}}\right)$ $\widetilde{O}\left(\frac{1}{\gamma^{2}T}\right)$ $\widetilde{O}\left(\frac{1}{\gamma^{2}\sqrt{n}}\right)$ Shallow
lei2024optimization ReLU $\widetilde{\Omega}\left(\frac{1}{\gamma^{8}}\right)$ $\widetilde{O}\left(\frac{1}{\gamma^{2}T}\right)$ $\widetilde{O}\left(\frac{1}{\gamma^{2}n}\right)$ Shallow
chen2021overparameterizationsufficientlearndeep ReLU $\widetilde{\Omega}\left(\frac{L^{22}}{\gamma^{8}}\right)$ $\widetilde{O}\left(\frac{1}{\gamma^{2}T}\right)$ $\widetilde{O}\left(\frac{e^{O(L)}}{\gamma}\sqrt{\frac{m}{n}}\right)$ Deep
taheri2025sharper Smooth $\widetilde{\Omega}\left(\frac{1}{\gamma^{6L+4}}\right)$ $\widetilde{O}\left(\frac{1}{\gamma^{2}T}\right)$ $\widetilde{O}\left(\frac{e^{O(L)}}{\gamma^{2}n}\right)$ Deep
@@
    We prove that gradient decent with step size $\eta$ and $T$ iterations achieves the convergence rate of $F_{S}(\overline{\mathbf{W}})/(\eta T)$, where $F_{S}(\overline{\mathbf{W}})=3\eta T\mathcal{L}_{S}(\overline{\mathbf{W}})+\|\overline{\mathbf{W}}-\mathbf{W}(0)\|_{F}^{2}$, $\overline{\mathbf{W}}$ is a reference model, $\mathbf{W}(0)$ is the initialization point and $\mathcal{L}_{S}(\cdot)$ is the training error. This indicates that all iterates remain within a neighborhood of $\overline{\mathbf{W}}$. By refining the analysis of ReLU activation patterns, we reduce the overparameterization requirement by a factor of $L^{6}$ as compared to chen2021overparameterizationsufficientlearndeep (see Remark 1 ).
@@
    We establish a population risk bound of $\widetilde{O}(L^{4}F(\overline{\mathbf{W}})/n)$, where $F(\overline{\mathbf{W}})$ extends the empirical counterpart $F_{S}(\overline{\mathbf{W}})$ to the population setting. Our analysis introduces two key technical contributions. First, we derive sharper Rademacher complexity bounds for the hypothesis class encompassing all gradient descent iterates. A central innovation is our representation of the complexity via products of sparse matrices, whose norms are tightly controlled using optimization-informed estimates (see Remark 2). Second, by leveraging the covering number techniques, we prove that ReLU networks remain $\widetilde{O}(L^{2})$-Lipschitz continuous in a neighborhood around the initialization—a substantial improvement over previous exponential bounds (JMLR:v25:23-0740; taheri2025sharper) (see Remark 3).
@@
    For NTK separable data with a margin $\gamma$, we show that neural networks with $m=\widetilde{O}(L^{16}\log({n/\delta})(\log n)^{8}/\gamma^{8})$ can achieve $\widetilde{O}(L^{4}(1+\gamma L^{2})/(\gamma^{2}n))$ risk. This improves on chen2021overparameterizationsufficientlearndeep’s $\widetilde{O}(e^{O(L)}\gamma^{-1}\sqrt{m/n})$ by simultaneously (a) removing the exponential depth dependence, (b) eliminating the $\sqrt{m}$ width factor and (c) achieving the optimal dependence on sample complexity (see Table 1 for a comparison with existing work).
@@ 2.1 Optimization
The foundational work of jacot2018neural introduced the Neural Tangent Kernel (NTK) framework, demonstrating that in the infinite-width limit, neural networks behave as linear models with a fixed tangent kernel (NEURIPS2020_b7ae8fec; lee2019wide). This lazy training regime (chizat2019lazy), where parameters remain close to initialization, enables gradient descent to converge to global optima near initialization (pmlr-v97-du19c; pmlr-v97-arora19a). These analyses showed that the training dynamics can be governed by the NTK Gram matrix, which leads to substantial overparameterization ($m\gtrsim n^{6}/\lambda_{0}^{4}$). This was later improved by oymak2020toward. They showed that if the square-root of the number of the network parameters exceeds the size of the training data, randomly initialized gradient descent converges at a geometric rate to a nearby global optima. The work (ji2020polylogarithmicwidthsufficesgradient) achieved polylogarithmic width requirements for logistic loss by leveraging the 1-homogeneity of two-layer ReLU networks. However, it should be mentioned that this special property does not hold for deep networks. The NTK framework was extended to deep architectures by arora2019exact for CNNs and by du2019gradientdescentprovablyoptimizes for ResNets using the last-layer NTK. JMLR:v25:23-0740 pointed out that such a characterization is loose, only capturing the contribution from the last layer. They further gave the uniform convergence of all layers as $m\to\infty$ and convergence guarantee for stochastic gradient descent (SGD) in streaming data setting. pmlr-v97-allen-zhu19a showed that the optimization landscape is almost-convex and semi-smooth, based on which they proved that SGD can find global minima. NEURIPS2019_cf9dc5e4 introduced the neural tangent random feature and showed the convergence of SGD under the overparameterized assumption $m\gtrsim n^{7}$.
@@ 2.2 Generalization
The NTK framework has yielded generalization bounds scaling as $\sqrt{\mathbf{y}^{\top}(\mathbf{H}^{\infty})^{-1}\mathbf{y}/n}$ (pmlr-v97-arora19a; NEURIPS2019_cf9dc5e4). This data-dependent complexity measure helps to distinguish between random labels and true labels. li2018learning showed that SGD trained networks can achieve small test error on specific structured data. A very popular approach to studying the generalization of neural networks is via the uniform convergence, which analyzes generalization gaps in a hypothesis space using tools such as Rademacher complexity or covering numbers (neyshabur2015norm; bartlett2017spectrally; golowich2018size; JMLR:v25:22-1250). However, this could lead to vacuous generalization bound in some cases (nagarajan2019uniform). Moreover, these bounds typically exhibit exponential dependence on depth $L$, thus often leading to loose bounds (chen2021overparameterizationsufficientlearndeep). This capacity-based method usually results in the generalization rate of the order $\widetilde{O}(1/\sqrt{n})$. Recent work has also exploited stability arguments for generalization guarantees (richards2021stability; NEURIPS2022_fb8fe6b7; JMLR:v25:23-0422; deora2023optimization; taheri2025sharper). The main idea of algorithmic stability is to study how the perturbation of training samples would affect the output of an algorithm (rogers1978finite). The connection to generalization bound was established in bousquet2002stability. hardt2016train gave the stability analysis of SGD for Lipschitz, smooth and convex problems. lei2020fine further studied SGD under much wilder assumptions. NEURIPS2020_b7ae8fec identified weak convexity of neural networks, enabling stability analyses with polynomial width requirements for quadratic loss (richards2021stability; NEURIPS2022_fb8fe6b7). Moreover, JMLR:v25:23-0422; taheri2025sharper obtained generalization bounds of order $\widetilde{O}(1/n)$ by using a generalized local quasi-convexity property for sufficiently parameterized networks. However, these methods depend on smooth activations, and whether similar or even better bound can be established for deep ReLU networks is still unknown. The recent work derived excess risk bounds of order $\widetilde{O}(1/n)$ for shallow ReLU networks (lei2024optimization).
@@ 3 Preliminaries
where we use logistic loss $\ell(z):=\log(1+\exp(-z))$ throughout this paper. We further assume the following symmetric initialization (nitanda2020optimal; kuzborskij2023learninglipschitzfunctionsgdtrained; JMLR:v25:23-0740):
@@
We remark that this initialization is for theoretical simplicity, using general initialization techniques will not affect the main results. We fix the output weights $\{a_{r}\}$ and use Gradient Descent (GD) to train the weight matrix $\mathbf{W}$ (ji2020polylogarithmicwidthsufficesgradient; pmlr-v97-arora19a; zou2018stochasticgradientdescentoptimizes).
@@
Note that in each layer we employ $\sqrt{{2}/{m}}$ as the regularization factor instead of the conventional $\sqrt{{1}/{m}}$ (ji2020polylogarithmicwidthsufficesgradient), which is due to $\mathbb{E}_{x\sim\mathcal{N}(0,1)}\sigma^{2}(x)={1}/{2}$ for our activation function $\sigma(\cdot)$. This scaling matches both the theoretical framework of pmlr-v97-du19c and the initialization scheme of (he2015delving) (where weights $\mathbf{w}_{r}^{l}\sim\mathcal{N}(0,{2}/{m})$). As will be shown later (Appendix A), this regularization ensures stable gradient propagation and maintains consistent variance across layers.
The following assumption is standard in the literature (NEURIPS2019_cf9dc5e4; ji2020polylogarithmicwidthsufficesgradient; chen2021overparameterizationsufficientlearndeep).
@@
where the first term captures the generalization error, quantifying the network’s performance on unseen data. The second term represents the optimization error, which reflects GD’s ability to find global minima. We will use tools in the optimization theory to study the optimization error (ji2020polylogarithmicwidthsufficesgradient; schliserman2022stability), and Rademacher complexity to control the generalization error (mohri2018foundations).
@@ 4.1 Optimization Analysis
Our theorem shows that the convergence rate is bounded by the optimization error of a reference model, implying that any low-loss reference point guarantees good convergence. While prior works relied on NTK-induced solutions (richards2021stability; pmlr-v97-arora19a), we prove that there exists a reference model near initialization under the milder Assumption 3. Furthermore, our analysis implies that all training iterates remain within a neighborhood of the reference point, and thus near initialization, aligned with NTK-based observations but without studying the kernel or the corresponding Gram matrix (pmlr-v97-du19c; du2019gradientdescentprovablyoptimizes).
@@
Telescoping gives the theorem. chen2021overparameterizationsufficientlearndeep introduce the following neural tangent random feature (NTRF) function class: